a train starting from a railway station and moving with uniform acceleration attains a speed of 40km/h in 10 minutes find its aaceleration
1
Expert's answer
2020-09-15T08:19:47-0400

As per the given question,

Initial velocity of train "(u)=0"

Final speed of the train "(v)=40km/hour = \dfrac{40\times 1000m}{3600sec}=\dfrac{100 m}{9sec}"

Time "(t)=10min = 10\times 60sec = 600sec"

Let the acceleration of the train is "=a"

We know that "v=u+at"

Now, substituting the values,

"\dfrac{100m}{9sec}=0+a\times 600 sec"

"\Rightarrow a = \dfrac{100m}{9\times 600sec^2}"

"\Rightarrow a = \dfrac{1}{18}m/sec^2"

"\Rightarrow a = 0.056m/sec^2"